OpenStudy (anonymous):

A taxi service has two rates, as follows: Rate A charges a flat $2.50 for every mile; rate B charges $15 plus $1.50 per mile. At least how many miles (expressed as a whole number) must one travel to make Rate B cheaper? A. 4 B. 15 C. 16 D. 17

OpenStudy (anonymous):

Can anyone please help and explain how to solve this?

OpenStudy (anonymous):

to make rate B cheaper you have to make rate A more expensive, the answer is 15 because $15.00 diveded by $2.50 equals 6, but then there is the charge of$1.50 in rate B so $1.50 times 6 is 9. 9+6 +15. hope this helped!

OpenStudy (anonymous):

actually make that 16

OpenStudy (anonymous):

because 15 makes it even 16 makes rate B cheaper

OpenStudy (anonymous):

in my first answer i meant 9+6=15

OpenStudy (anonymous):

In the first answer, is it 15 or 16?

OpenStudy (anonymous):

Because I am confused:(

OpenStudy (anonymous):

16 :)

OpenStudy (anonymous):

So it's to make rate B cheaper you have to make rate A more expensive, the answer is 16 because $16.00 diveded by $2.50 equals 6, but then there is the charge of$1.50 in rate B so $1.50 times 6 is 9. 9+7 =16.

OpenStudy (anonymous):

Is this how it should be?

OpenStudy (anonymous):

If you don't mind can you explain from the beginning because I am confused:(

OpenStudy (anonymous):

Ill retype it hopefully to make sense lol ok so, To make rate B cheaper you need to make rate A more expensive, the answer is 16 because $15.00 (from rate B) divided by $2.50 (from rate A) equals 6 miles, but then there is the charge of $1.50 per mile in rate B so $1.50 x 6 (the miles used in rate A) equals 9 miles, 9+6=15 miles but to make Rate B have more miles you add one more, so it would be 16 miles. That would make Rate A paying $40 and rate B paying $39.

OpenStudy (anonymous):

I meant (minor error), To make rate B cheaper you need to make rate A more expensive, the answer is 16 because $15.00 (from rate B) divided by $2.50 (from rate A) equals 6, but then there is the charge of $1.50 per mile in rate B so $1.50 x 6 (the miles used in rate A) equals 9 miles, 9+6=15 miles but to make Rate B have more miles you add one more, so it would be 16 miles. That would make Rate A paying $40 and rate B paying $39.

OpenStudy (anonymous):

lol I meant this!! To make rate B cheaper you need to make rate A more expensive, the answer is 16 because $15.00 (from rate B) divided by $2.50 (from rate A) equals 6, but then there is the charge of $1.50 per mile in rate B so $1.50 x 6 (the miles used in rate A) equals 9 miles, 9+6=15 miles but to make Rate B cheaper you add one more miles, so it would be 16 miles. That would make Rate A paying $40 and rate B paying $39.

OpenStudy (anonymous):

@LeaRD so the third one is the correct one right?!

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

Thaaanks :D